# Artin–Schreier quandles of involutions in absolute Galois groups

### 12 Mar 2024 v1

@@
Artin–Schreier quandles of involutions in absolute Galois groups
@@
We introduce a new invariant of fields that refines their real spectrum and is related to their absolute Galois group: the Artin–Schreier quandle. For formally real number fields, it is freely generated in its variety by a Cantor space of indeterminates. For Laurent series fields, we compute it in terms of the Artin–Schreier quandle of the coefficient field. This result and other examples show that, in general, there are relations.
@@ 1 Introduction
Recall that we can identify the real spectrum of a field $F$ with the set of conjugacy classes of involutions in the absolute Galois group $\operatorname{Gal}(F)$. Here, instead of working with equivalence classes, we define the Artin–Schreier quandle $\mathrm{AS}(F)$ directly, building on the set of all involutions inside the absolute Galois group $\operatorname{Gal}(F)$. Novel is the observation that this set not only has a topological structure as a subspace of a pro-finite group, but it also has an algebraic structure. While not a subgroup, it is still closed under conjugation. The algebraic theory of quandles is a well-established axiomatic framework to model conjugation, and $\mathrm{AS}(F)$, therefore, has the structure of an involutory quandle. One of our main results describes the pro-finite involutory quandle $\mathrm{AS}(F)$ in the most fundamental cases:
@@
The Artin–Schreier quandle $\mathrm{AS}(\mathbb{Q})$ of the rational number field $\mathbb{Q}$ is a free pro-finite involutory quandle. A Cantor space of involutions inside the absolute Galois group $\operatorname{Gal}(\mathbb{Q})$ gives a basis. A similar statement holds for all formally real number fields.
We prove this result as Theorems 4.1 and 4.6 in the text. Totally imaginary number fields have no involutions in their absolute Galois groups, and their Artin–Schreier quandles are empty.
For context, recall that Hilbert’s 17th problem asked if we could write a real multivariate polynomial that only assumes non-negative values as a sum of squares of rational functions. This problem was solved affirmatively in 1927 by Artin [2] who, in collaboration with Schreier [1, 3, 4], founded the subject of real algebra. In 1979, Coste and Roy introduced the real spectrum of a commutative ring [13]. It plays the same role in real algebraic geometry as the ordinary prime spectrum in the usual algebraic geometry. For a field $F$, the real spectrum has various, seemingly different descriptions: If $F$ is a number field, we can identify it with the finite set of embeddings of $F$ into the field $\mathbb{R}$ of real numbers. In general, it is better to think of it as the set of orderings of $F$, as the set of ring morphisms $\mathrm{W}(F)\to\mathbb{Z}$, where $\mathrm{W}(F)$ is the Witt ring of quadratic forms over $F$, or—for us—as the set of conjugacy classes of involutions in the absolute Galois group $\operatorname{Gal}(F)$. The Artin–Schreier quandle replaces the set of conjugacy classes with the collection of all their representatives and non-judgementally embraces the algebraic structure that comes with it.
The thrust of this idea extends beyond real algebra to the many other contexts of conjugation in algebraic number theory, and it opens new avenues in the flourishing ‘arithmetic topology’ program initiated by Manin, Mazur, Mumford, and others. The connection to topology comes via the theory of quandles. We can think of these algebraic structures as generalisations of knots: first considered by Burde [12] in that context, Joyce and Matveev [29, 38], building on earlier work of Waldhausen, showed that knot quandles, a canonical refinement of knot groups, completely classify knots. The knot quandle of a prime knot is the conjugacy class of the meridian with the quandle structure, again, given by conjugation.
Meridians are generators of the fundamental groups of ‘small’ disks that are transversal to the knot and have been punctured by removing the knot. Given the importance of meridians for the understanding of knot groups and quandles, we are urged to understand the function fields $F(\!(t)\!)$ of punctured formal disks. The following result is Theorem 5.1 in the main text.
@@
$\mathrm{AS}(F(\!(t)\!))\cong\mathrm{AS}(F)\times\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$
of involutory quandles, where $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ is the free pro-finite involutory quandle on a set $\{\sigma,\tau\}$ with two elements.
This paper is written to be accessible to number theorists and topologists. We discuss involutions and involutory quandles in Section 2. The second half of that section describes the free models and covers the topological details needed for the pro-finite context. We specialise this general theory to Artin–Schreier quandles in Section 3. We describe the Artin–Schreier quandles of all number fields and other global fields in Section 4. The final Section 5 contains Theorem 5.1 and discusses other examples of Artin–Schreier quandles with relations.
@@ 2 Involutions and quandles
We start with some necessary recollections and extensions to the pro-finite setting of the main algebraic structure that we are interested in: quandles, which describe symmetries in a way different from, but of course related to, groups. Suitable references are [10, 17, 29, 38]. The terminology and notation varies considerably among the various sources, and we shall stick to the following throughout:
@@
for all all $x$$y$, and $z$ in $R$.
@@
By definition, racks bring their own automorphisms, namely the left-multiplications, one for each element $x$, but there is also a natural automorphism, namely $x\mapsto x\rhd x$, which is compatible with all quandle morphisms and which generates all natural automorphisms [48, Thm. 5.4], and which not necessarily has the form of a left-multiplication.
@@
for all elements $x$.
@@
The underlying set of any group forms a quandle under conjugation, i.e., if we define $g\rhd x=gxg^{-1}$. Any pro-finite group gives rise to a pro-finite quandle in this way.
@@
An involutory quandle is sometimes called a kei, following Takasaki [50]. The term Takasaki quandle is sometimes used in his honour. Our present terminology follows Joyce [29]. Fenn and Rourke [17, Sec. 1.3] call such structures involutive, but according to Brieskorn [10, Sec. 2], this term refers to a completely different property that we will not need here. We refer to [29, Sec. 10,11] for more information on involutory quandles.
@@
If $G$ is a group, the formula $g\rhd x=gx^{-1}g$ defines a binary operation that turns the whole underlying set of $G$ into an involutory quandle. This class of examples seems to originate in Loos’ work [35] on the structure of symmetric spaces. As a special case, if $G$ is the Euclidean plane, with the group structure given by addition, then we have $g\rhd x=2g-x$, so that left-multiplication is reflection in the point $g$. Figure 1 illustrates the rack axiom.
$x$$y$$z$$y\rhd z$$x\rhd z$$x\rhd y$$x\rhd(y\rhd z)=(x\rhd y)\rhd(x\rhd z)\phantom{nothinghere}$
@@
$\operatorname{Inv}(G)=\{s\in G\,|\,s^{2}=\operatorname{id}\not=s\}$ (2.4)
of involutions in $G$ forms an involutory quandle with respect to conjugation. We emphasise that this convention excludes the identity from being an involution. If $G$ is pro-finite, and $\operatorname{Inv}(G)$ is a closed subset, then $\operatorname{Inv}(G)$ will also be pro-finite. We shall see below, in Lemma 3.7, that this is always the case in the situations we are interested in.
Any algebraic theory in the sense of Lawvere [34] is determined by its free models. For the theory of groups, for instance, we have the free groups $\mathrm{F}(S)$ generated by sets $S$, and if $S$ has $n$ elements, then the elements of $\mathrm{F}(S)$ are in bijection with the natural operations on groups of arity $n$. Specifically, if $S=\{x,y\}$, then the multiplication corresponds to $xy\in\mathrm{F}\{x,y\}$, whereas $xyx^{-1}\in\mathrm{F}\{x,y\}$ gives conjugation. For the theory of quandles, we can describe the free models as follows.
@@
We refer to [29, Thm. 4.1] and [31, Satz 2.5] for proofs.
We can also describe free involutory quandles explicitly in a similar manner. We will have to replace the free groups $\mathrm{F}(S)$, which are the universal Artin groups, with the universal Coxeter groups: The universal Coxeter group on a set $S$ of generators is the quotient
$\mathrm{F}_{2}(S)=\langle\,s\in S\,|\,s^{2}=e\text{ for all }s\in S\,\rangle$
of the free group $\mathrm{F}(S)$ by the relations that say that the generating elements in $S$ are involutions. When $S$ has $n$ elements, then $\mathrm{F}_{2}(S)\cong(\mathbb{Z}/2)^{\star n}$ is the coproduct of $n$ copies of the group of order $2$.
@@
We refer to [29, Cor. 10.3] and [31, Satz 2.5] once more.
@@
We note that $\mathrm{F}_{2}(S)$ is not a free object in any variety of groups. The equation $s^{2}=e$ is required for the generators $s$ only, not necessarily for every group element. (The stronger condition would force the group to be abelian.) In contrast, the object $\operatorname{FQ}_{2}(S)$ is free in the theory of involutory quandles.
@@
We have $\operatorname{FQ}_{2}(S)=\{\,g\in\mathrm{F}_{2}(S)\,|\,g^{2}=e\not=g\,\}$.
@@
Clearly, every conjugate of a generator $s\in S$ in $\mathrm{F}_{2}(S)$ is an involution. The converse also holds: every involution is conjugate to a generator. This is elementary to verify, and it is also a consequence of the Kurosh subgroup theorem; see [37, Cor. 4.1.4], for instance. ∎
@@
We can rephrase the observation leading to Proposition 2.13 in more conceptual terms as follows (compare [7] and [8, Sec. 10.11]). In any category with finite sums, we can take the full subcategory spanned by the finite sums of any chosen object $Z$, and this will be an algebraic theory in the sense of Lawvere. For instance, if we consider the category of models for an algebraic theory, and choose for $Z$ a free model on one generator, we recover that algebraic theory in this way. In particular, for the group $Z=\mathbb{Z}$, the free groups recover the algebraic theory of groups. More generally, we now see that any group $Z$ defines an algebraic theory such that the sets $\operatorname{Mor}(Z,G)$ of group morphisms from $Z$ into groups $G$ are models for that theory. For the case $Z=\mathbb{Z}/2$ of interest to us here, we see that the sets $\operatorname{Mor}(\mathbb{Z}/2,G)\cong\operatorname{Inv}(G)\cup\{e\}$ support models for an algebraic theory: the theory of involutory quandles with unit, where a unit $u$ is an element that is both fixed and fixing: $x\rhd u=u$ and $u\rhd y=y$ for all $x,y$ (see [33, Def. 2.8], for example).
For the Galois theory context, we also need the pro-finite versions $\mathrm{F}^{\wedge}(S)$ and $\mathrm{F}^{\wedge}_{2}(S)$ of the groups $\mathrm{F}(S)$ and $\mathrm{F}_{2}(S)$, respectively, where $S$ can now be any pro-finite space. Suitable references come from Gildenhuys et al. [23, 24], which build on earlier work of Neukirch et al. [41, 42, 9], Mac Lane [36], and the writings of Haran and Jarden [25, 26], of course.
@@
If the space $S$ is finite, we can build $\mathrm{F}^{\wedge}(S)$ as the pro-finite completion of $\mathrm{F}(S)$, namely the limit $\lim_{N}\mathrm{F}(S)/N$, where $N$ ranges over the finite index normal subgroups of $\mathrm{F}(S)$. If $S$ is, more generally, pro-finite, we have two choices: on the one hand, we can set $\mathrm{F}^{\wedge}(S)=\lim_{N}\mathrm{F}(S)/N$ as before, but require that $gN\cap S$ is open in $S$ for the (finitely many) cosets $gN$ of $N$ in $\mathrm{F}(S)$. On the other hand, we can set $\mathrm{F}^{\wedge}(S)=\lim_{R}\mathrm{F}^{\wedge}(S/R)$, where $R$ ranges over the equivalence relations on $S$ that have a finite quotient. Both constructions give isomorphic results. For $\mathrm{F}^{\wedge}_{2}(S)$, we use $\mathrm{F}^{\wedge}_{2}(S)=\lim_{N}\mathrm{F}(S)/N$ with the added requirement that $N$ contains all squares of elements in $S$ or work with $\mathrm{F}^{\wedge}_{2}(S)=\lim_{R}\mathrm{F}^{\wedge}_{2}(S/R)$. Either way, a continuous homomorphism $\mathrm{F}_{2}(S)\to G$ into a pro-finite group $G$ is the same thing as a continuous map $S\to G$ such that the elements in the image square to the identity element $e$. There are pro-finite versions of the Kurosh subgroup theorem:
@@
Every involution in $\mathrm{F}^{\wedge}_{2}(S)$ is conjugate to an element in $S$.
@@
If $S$ is a finite set of $n$ elements, Prop. 6.1 in [25] shows that there are exactly $n$ conjugacy classes of involutions in $\mathrm{F}^{\wedge}_{2}(S)$, represented by the elements in $S$. In [26, Cor. 3.2], the authors extend this to the general case. ∎
@@
For every pro-finite space $S$, there is a pro-finite involutory quandle $\mathrm{FQ}^{\wedge}_{2}(S)$ together with a continuous map $S\to\mathrm{FQ}^{\wedge}_{2}(S)$ that satisfies the universal property: a continuous homomorphism $\mathrm{FQ}^{\wedge}_{2}(S)\to Q$ into a pro-finite involutory quandle $Q$ is, via restriction, the same thing as a continuous map $S\to Q$.

### 08 Apr 2026 v2

@@
Artin–Schreier quandles of involutions
in absolute Galois groups
@@
We introduce a new invariant of fields that refines their real spectrum and is related to their absolute Galois group: the Artin–Schreier quandle. For totally real number fields, it is freely generated—in the variety of profinite involutory quandles—by a Cantor space of indeterminates. For Laurent series fields, we compute it in terms of the Artin–Schreier quandle of the coefficient field. This result, along with other examples, shows that, in general, there are relations.
@@ 1 Introduction
The real spectrum of a commutative ring is a topological space that enriches its Zariski spectrum by incorporating ideas from the theory of formally real fields. For a field, which has only one prime ideal, the Zariski spectrum is merely a singleton, whereas the real spectrum comprises the space of orderings of the field. From a group-theoretic perspective, we know that we can identify the real spectrum of a field with the space of conjugacy classes of involutions in the absolute Galois group (see Section 3 for a review). This observation allows us, in this paper, to further refine the real spectrum of a field $F$. Instead of working with equivalence classes in the form of conjugacy classes of involutions, we define the Artin–Schreier quandle $\mathrm{AS}(F)$ directly, based on the set of all involutions within the absolute Galois group $\operatorname{Gal}(F)=\operatorname{Gal}(\overline{F}|F)$. A novel observation is that this set not only possesses a topological structure as a subspace of a profinite group but also an algebraic one: Although it is not a subgroup, it is still closed under conjugation. The algebraic theory of quandles is a well-established axiomatic framework for modelling conjugation, and $\mathrm{AS}(F)$, therefore, inherits the structure of an involutory quandle (see Section 2 for a review). One of our main results characterises the profinite involutory quandle $\mathrm{AS}(F)$ in the most fundamental cases:
@@
The Artin–Schreier quandle $\mathrm{AS}(\mathbb{Q})$ of the rational number field $\mathbb{Q}$ is a free profinite involutory quandle. A Cantor space of involutions inside the absolute Galois group $\operatorname{Gal}(\mathbb{Q})$ gives a basis. A similar statement holds for all totally real number fields.
We prove this result as Theorems 4.1 and 4.4 in the text. Totally imaginary number fields have no involutions in their absolute Galois groups, and their Artin–Schreier quandles are empty.
For more context, recall that Hilbert’s 17th problem asked if we could write a real multivariate polynomial that only assumes non-negative values as a sum of squares of rational functions. This problem was solved affirmatively in 1927 by Artin [2] who, in collaboration with Schreier [1, 3, 4], founded the subject of real algebra. In 1979, Coste and Roy introduced the real spectrum of a commutative ring [13]. It plays the same role in real algebraic geometry as the ordinary prime spectrum in the usual algebraic geometry. For a field $F$, the real spectrum has various, seemingly different descriptions: If $F$ is a number field, we can identify it with the (finite) set of embeddings of $F$ into the field $\mathbb{R}$ of real numbers. In general, referring again to the review in Section 3 for more detail, it is better to think of it as the set of orderings of $F$, as the set of ring morphisms $\mathrm{W}(F)\to\mathbb{Z}$, where $\mathrm{W}(F)$ is the Witt ring of quadratic forms over $F$, or—for us—as the set of conjugacy classes of involutions in the absolute Galois group $\operatorname{Gal}(F)$. The Artin–Schreier quandle replaces the set of conjugacy classes with the collection of all their representatives and non-judgementally embraces the algebraic structure that comes with it.
The thrust of this idea extends beyond real algebra to the many other contexts of conjugation in algebraic number theory, and it opens new avenues in the flourishing ‘arithmetic topology’ program initiated by Manin, Mazur, Mumford, and others. The connection to topology comes via an algebraic structure that codifies conjugation: quandles. We refer to Section 2 again for a review of the basic definition, but we point out here that we can think of quandles as generalisations of knots: first considered by Burde [12] in that context, Joyce [30] and Matveev [42], building on earlier work of Waldhausen, showed that knot quandles, a canonical refinement of knot groups, completely classify knots. The knot quandle of a prime knot is the conjugacy class of the meridian with the quandle structure given by conjugation. (For composite knots, the description is more subtle and not relevant here.) Meridians are generators of the fundamental groups of ‘small’ disks that are transversal to the knot and have been punctured by removing the intersection with the knot. Given the importance of meridians for understanding knot groups and quandles, we are urged to understand the function fields F((t))F(\!(t)\!) of punctured formal disks. The following result is Theorem 5.1 in the main text.
@@
$\mathrm{AS}(F(\!(t)\!))\cong\mathrm{AS}(F)\times\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}\cong\mathrm{AS}(F)\times\widehat{\mathbb{Z}}$
of profinite involutory quandles, where $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ is the free profinite involutory quandle on a set $\{\sigma,\tau\}$ with two elements, and which, as a profinite space, is homeomorphic to $\widehat{\mathbb{Z}}$. The quandle structure is given by $a\rhd b=2a-b$ for $a,b\in\widehat{\mathbb{Z}}$.
We note that Takahashi [57] (see also Davis–Schlank [14]) has associated quandles with Galois covers of arithmetic schemes in a different way than here. We refer to Shusterman [53] for another framework that uses quandles and the more general racks in number theory.
This paper is written to be broadly accessible to number theorists and topologists. We discuss involutions and involutory quandles in Section 2, with the second half of that section describing the free models and covering the topological details needed for the profinite context. We specialise this general theory to Artin–Schreier quandles in Section 3. We describe the Artin–Schreier quandles of number fields and other global fields in Section 4. The final Section 5 contains Theorem 5.1 and discusses examples of Artin–Schreier quandles that have relations.
@@ 2 Involutions and quandles
We start with some necessary recollections and extensions to the profinite setting of the main algebraic structure that we are interested in: quandles, which describe symmetries in a way different from, but of course related to, groups. Suitable references are [10, 17, 30, 42]. The terminology and notation vary considerably among the various sources, and we shall stick to the following throughout:
@@
for all xxyy, and $z$ in $R$. A profinite rack is a profinite space $R$ together with a rack structure $\rhd\colon R\times R\to R$ that is continuous.
Recall that a profinite space is a totally disconnected compact Hausdorff space, where a compact Hausdorff space is totally disconnected if we can separate any two points by subsets that are both open and closed at the same time.
@@
By definition, racks bring their own automorphisms, namely the left-multiplications, one for each element $x$. Still, there is also a natural automorphism, namely $x\mapsto x\rhd x$, which is compatible with all rack morphisms and which generates all natural automorphisms [55, Thm. 5.4], and which does not necessarily have the form of a left-multiplication.
@@
for all elements $x$. A profinite quandle is a profinite rack that is a quandle.
@@
The underlying set of any group forms a quandle under conjugation, i.e., if we define $g\rhd x=gxg^{-1}$. Any profinite group gives rise to a profinite quandle in this way.
@@
An involutory quandle is sometimes called a kei, following Takasaki [58]. The term Takasaki quandle is sometimes used in his honour. Our present terminology follows Joyce [30]. Fenn and Rourke [17, Sec. 1.3] call such structures involutive, but according to Brieskorn [10, Sec. 2], this term refers to a completely different property that we will not need here. We refer to [30, Sec. 10f.] for more information on involutory quandles.
@@
If $G$ is a group, the formula $g\rhd x=gx^{-1}g$ defines a binary operation that turns the whole underlying set of $G$ into an involutory quandle. This class of examples seems to originate in Loos’ work [39] on the structure of symmetric spaces. As a particular case, if $G$ is the Euclidean plane, with the group structure given by addition, then we have $g\rhd x=2g-x$, so that left-multiplication is reflection in the point $g$. Figure 1 illustrates the rack axiom.
[…]
@@
$\operatorname{Inv}(G)=\{s\in G\mid s^{2}=e\not=s\}$ (2.4)
of involutions in $G$ forms an involutory quandle with respect to conjugation. We emphasise that this convention excludes the identity from being an involution. If $G$ is profinite, and $\operatorname{Inv}(G)$ is a closed subset, then $\operatorname{Inv}(G)$ will also be profinite. We shall see below, in Lemma 3.7, that this is always the case in the situations we are interested in.
To understand an algebraic structure, it is vital to understand its free models. In fact, any algebraic theory is determined by its free models (see Lawvere [38]), but we will not use the language of algebraic theories here except for in Remark 2.14 further down. However, we do need to discuss the free models in the relevant examples. For the theory of groups, for instance, we have the free groups $\mathrm{F}(S)$ generated by sets $S$, and if $S$ has $n$ elements, then the elements of $\mathrm{F}(S)$ are in bijection with the natural operations on groups of arity $n$. Specifically, if $S=\{x,y\}$, then the multiplication corresponds to the element $xy\in\mathrm{F}\{x,y\}$, whereas $xyx^{-1}\in\mathrm{F}\{x,y\}$ gives conjugation. For the theory of quandles, we can describe the free models as follows.
@@
We refer to [30, Thm. 4.1] and [34, Satz 2.5] for proofs.
We can also describe free involutory quandles explicitly in a similar manner. We will have to replace the free groups $\mathrm{F}(S)$, which we can think of as the universal Artin groups, with the universal Coxeter groups. Recall that an Artin group can be presented in the form $\langle S|R\rangle$, where $S$ is a set of generators, and $R$ is a set of relations of the form sts=tststs...=tst... for $s\not=t$ in $S$. The universal Artin group corresponds to the case $R=\text{\O }$, i.e., just the free group generated by the set $S$. Coxeter groups have the additional relations $s^{2}=e$ for all generators $s$ in $S$, so that the universal Coxeter group [45] on a set $S$ of generators is the quotient
$\mathrm{F}_{2}(S)=\langle\,s\in S\mid s^{2}=e\text{ for all }s\in S\,\rangle$
of the free group $\mathrm{F}(S)$ by the relations that say that the generating elements in $S$ are involutions. When the set $S$ has $n$ elements, then the group $\mathrm{F}_{2}(S)\cong(\mathbb{Z}/2)^{\star n}$ is the coproduct of $n$ copies of the group of order $2$.
@@
We refer to [30, Cor. 10.3] and [34, Satz 2.5] once more.
@@
We remark that the group $\mathrm{F}_{2}(S)$ is not defined by universal identities or laws; the equation $s^{2}=e$ is required for the generators $s$ only. It does not necessarily hold for every group element. If we require $g^{2}=e$ for all elements $g$ in a group $G$, the group $G$ must be abelian. In fact, this law defines the variety of elementary abelian 22–groups, and $\mathrm{F}_{2}(S)$ is not abelian unless $S$ has at most one element. So, while the functor $\mathrm{F}_{2}$ from the category of sets to the category of groups is a left-adjoint, it is not left-adjoint to the canonical forgetful functor. In contrast, the involutory quandle $\operatorname{FQ}_{2}(S)$ is characterised as a left-adjoint to the forgetful functor from the category of involutory quandles; it is a free object among involutory quandles.
@@
We have $\operatorname{FQ}_{2}(S)=\{\,g\in\mathrm{F}_{2}(S)\mid g^{2}=e\not=g\,\}$.
@@
Clearly, every conjugate of a generator $s\in S$ in $\mathrm{F}_{2}(S)$ is an involution. The converse also holds: Every involution is conjugate to a generator. This is elementary to verify, and it is also a consequence of the Kurosh subgroup theorem; see [41, Cor. 4.1.4], for instance. ∎
@@
We can rephrase the observation leading to Proposition 2.13 in more conceptual terms as follows. In any category with finite sums, we can take the full subcategory whose objects are the finite sums of any chosen object $Z$, and this will be an algebraic theory in the sense of Lawvere [38]. For instance, if we consider the category of models for an algebraic theory and choose for $Z$ a free model on one generator, we recover that algebraic theory in this way. In particular, for the group $Z=\mathbb{Z}$, the free groups recover the algebraic theory of groups. More generally, we now see that any group $Z$ defines an algebraic theory such that the sets $\operatorname{Mor}(Z,G)$ of group morphisms from $Z$ into groups $G$ are models for that theory. For the case $Z=\mathbb{Z}/2$ of interest to us here, we see that the sets $\operatorname{Mor}(\mathbb{Z}/2,G)\cong\operatorname{Inv}(G)\cup\{e\}$ support models for an algebraic theory: the theory of involutory quandles with unit, where a unit $u$ is an element that is both fixed and fixing: $x\rhd u=u$ and $u\rhd y=y$ for all $x,y$ (see [37, Def. 2.8] for this terminology). In fact, the theory defined by $Z=\mathbb{Z}/2$ is the theory of involutory quandles with unit. This insight goes back to Bergman (see his ICM talk [7] for general statements and his textbook [8, Sec. 10.11] for details).
@@
For the Galois theory context, we also need the profinite versions $\mathrm{F}^{\wedge}(S)$ and $\mathrm{F}^{\wedge}_{2}(S)$ of the groups $\mathrm{F}(S)$ and $\mathrm{F}_{2}(S)$, respectively, where $S$ can now be any profinite space. If the space $S$ is finite, we can build the profinite groups $\mathrm{F}^{\wedge}(S)$ as the profinite completions of the free groups $\mathrm{F}(S)$, namely the limit $\lim_{N}\mathrm{F}(S)/N$, where $N$ ranges over the finite index normal subgroups of $\mathrm{F}(S)$. If $S$ is, more generally, profinite, we have two choices: On the one hand, we can set $\mathrm{F}^{\wedge}(S)=\lim_{N}\mathrm{F}(S)/N$ as before, but require that $gN\cap S$ is open in $S$ for the (finitely many) cosets $gN$ of $N$ in $\mathrm{F}(S)$. On the other hand, we can set $\mathrm{F}^{\wedge}(S)=\lim_{R}\mathrm{F}^{\wedge}(S/R)$, where $R$ ranges over the equivalence relations on $S$ that have a finite discrete quotient. Both constructions give isomorphic results [24, Prop. 1.7]. This and other suitable references come from Gildenhuys et al. [24, 25], which build on earlier work of Neukirch et al. [46, 47, 9], Mac Lane [40], and the writings of Haran and Jarden [26, 27]. The last two are especially relevant because of the next result, Proposition 2.15, which they provide. They concern the groups $\mathrm{F}^{\wedge}_{2}(S)$ that are characterised by the property that a continuous homomorphism $\mathrm{F}^{\wedge}_{2}(S)\to G$ into a profinite group $G$ is the same thing as a continuous map $S\to G$ such that the elements in the image square to the identity element $e$. To construct them, we can follow [27, §1 and Prop. 3.1] and use $\mathrm{F}^{\wedge}_{2}(S)=\lim_{N}\mathrm{F}(S)/N$ with the added requirement that $N$ contains all squares of elements in $S$ or work with $\mathrm{F}^{\wedge}_{2}(S)=\lim_{R}\mathrm{F}^{\wedge}_{2}(S/R)$. There are profinite versions of the Kurosh subgroup theorem:
@@
Every involution in $\mathrm{F}^{\wedge}_{2}(S)$ is conjugate to a unique element in $S$, and the centraliser of any involution consists of only that involution and the identity.
@@
If $S$ is a finite set of $n$ elements, Haran–Jarden [26, Prop. 6.1] show that there are exactly $n$ conjugacy classes of involutions in $\mathrm{F}^{\wedge}_{2}(S)$, represented by the elements in $S$. In [27, Cor. 3.2], they extend this to the general case. ∎
@@
For every profinite space $S$, there is a profinite involutory quandle $\mathrm{FQ}^{\wedge}_{2}(S)$ together with a continuous map $S\to\mathrm{FQ}^{\wedge}_{2}(S)$ that satisfies the universal property: Restriction gives a bijection between the set of continuous homomorphism $\mathrm{FQ}^{\wedge}_{2}(S)\to Q$ into a profinite involutory quandle $Q$ and the set of continuous maps $S\to Q$. Furthermore, there is a continuous map $\mathrm{FQ}^{\wedge}_{2}(S)\to S$ that identifies the target $S$ with the orbit space of the $\mathrm{F}^{\wedge}_{2}(S)$–action on $\mathrm{FQ}^{\wedge}_{2}(S)$ via conjugation.